# LSCodec: Low-Bitrate and Speaker-Decoupled Discrete Speech Codec thanks:

### v2

@@ III-D Stage-wise Evaluations and Ablation Study
Ablation on SSL token prediction task
No $\mathcal{L}_{\text{idx}}$ 11.22 0.811 8.93 0.816
Ablation on multi-stage training
No stage 1 training - - 3.84 0.800
IV Conclusion
We propose LSCodec, a low-bitrate and speaker-decoupled speech codec. LSCodec constructs a three-stage training framework with speaker perturbation. A VQ layer is applied after a VAE that disentangles speaker in a continuous space. A token vocoder is then trained upon the quantized codes. Experiments prove that LSCodec achieves strong reconstruction and VC ability given a bitrate as low as 0.25kbps. Ablation studies validate the proposed training framework. LSCodec paves the way of future LM-based speech generation for its highly compact representation space. Explorations on stronger perturbation methods and better prosody preservation are valuable for future work.

### v3

@@ 3.5 Ablation Study
On SSL token prediction No $\mathcal{L}_{\text{idx}}$ 11.22 0.811 8.93 0.816
On multi-stage training No stage 1 training - - 3.84 0.800
4 Conclusion
We propose LSCodec, a low-bitrate and speaker-decoupled speech codec with a highly compact representation space. LSCodec constructs a multi-stage training framework with speaker perturbation. A single-codebook VQ layer is applied after a VAE that disentangles speaker in a continuous space. A token vocoder is then trained upon the quantized codes. Experiments prove that LSCodec achieves strong reconstruction and VC ability given a bitrate as low as 0.45 and 0.25kbps. Probing experiments and ablation studies validate the proposed methods. Meanwhile, stronger perturbation methods, better content preservation, and scaling ability can be explored as future work.
5 Acknowledgements
This work is supported by China NSFC Project (No. 92370206), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), and Key Research and Development Program of Jiangsu Province, China (No.BE2022059).